OpenStudy (anonymous):

the temperature is given by T(H)=-A-Bcos(piH/12), T is temperature and H is the number of hours from midnight (0

OpenStudy (anonymous):

in your formula is Capital H and lower h supposed to be the same variable?

OpenStudy (anonymous):

oops yeah

OpenStudy (anonymous):

Just checking.

OpenStudy (anonymous):

Another thing the portion you have that says ( 0 <= T < 24). Are you sure that T doesn't say H?

OpenStudy (anonymous):

Sorry, this may sound like a stupid question that part really doesn't make sense.

OpenStudy (anonymous):

yup it's T, but now that you mentioned it it doesnt make sense

OpenStudy (anonymous):

Yea I believe that T should be an H, because H is time and if we consider military time, which has 24 hours in a day, then that inequality makes sense... hmmm

OpenStudy (anonymous):

i don't knowww, some one making this assignment made a mistake. let's go with it's supposed to be H

OpenStudy (anonymous):

Is we go with the inequality as (0 <= H < 24)... A.) -15 = -A - B*cos(0) --> -15 = -A - B 5 = -A - B*cos(pi) --> 5 = -A + B. you have two equations with two unknowns, all thats left is to solve.

OpenStudy (anonymous):

how did you go from -15 to 5.

OpenStudy (anonymous):

or did you solve for both?

OpenStudy (anonymous):

I solved two separate times The told us that when T= -15 H =0, and T=5 and H = 12

OpenStudy (anonymous):

and did you have to plug the 12 in for when you plugged the 5 degrees in cuz its T(H)?

OpenStudy (anonymous):

T(H) is just telling you that t is a function of H. Sorry i assumed to much left me redo what i did early to be more explicit with what I mean.

OpenStudy (anonymous):

ohh ok i get it. I was thinking that but for other things i've been going over i had to substitute in all the variables.

OpenStudy (anonymous):

We are told that initially (at midnight) the temperature is -15 degrees right? well... T(0) = -A - B*cos (pi*0/12) --> T(0) = -A - B*cos (0) -- >where T(0) = -15 Therefore, -15 = -A - B. Then i did same with other information to the same equation again and got 5 = -A + B understand?

OpenStudy (anonymous):

From there we have two equations and two unknowns.solve and tell me what you get and I'll tell you what i got.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got a=5 b=7

OpenStudy (anonymous):

Check you B again. show me your math.

OpenStudy (anonymous):

ohh b=10

OpenStudy (anonymous):

YUP =)

OpenStudy (anonymous):

Now part B....

OpenStudy (anonymous):

yeahh... i think it has something to do with something similair to the trapizoid method..

OpenStudy (anonymous):

Maybe, but we can do it another way. We can find all of the temperatures for the first ten hours then average them. So.... Are new equation is now T(H) =-5-10*cos(pi*H/12), since we found A and B. T(0) = ? T(1) = ? T(2) = ? . . . T(10) = ? from here [T(0) + T(1) + T(2) + ... + T(10)]/10 Does that make sense?

OpenStudy (anonymous):

To be honest I don't remember the trapezoidal method. Not much help form here on

OpenStudy (anonymous):

i think so.

OpenStudy (anonymous):

haha me eitherr, learned it like three months ago but nothing sticks with migraines 24/7

OpenStudy (anonymous):

Hmm. Well for part b linear interpolation should get you the right answer i believe. From here on I don't remember, I thought i did but my answers are coming up super wrong, not even close haha sorry.

OpenStudy (anonymous):

its all good, thanks for the help

OpenStudy (kropot72):

T(H)=-A-B\[T(H)=-A-B \cos (\frac{\pi H}{12})\] At midnight: \[T(H=0)=-A-B \cos( 0)=-A-(B \times 1)=-A-B=-15\] At noon: \[T(H=12)=-A-B \cos( \pi)=-A-(B \times -1)=-A+B=5\] -A -B = -15 -A + B = 5 Adding gives: -2A = -10 A = 5 By substitution: B = 10

OpenStudy (anonymous):

yepp

OpenStudy (kropot72):

The average temperature for the first 10 hours is found by integrating the formula between the limits H=0 and H=10 and then dividing the result by 10 to find the average. Do you follow?

OpenStudy (kropot72):

\[\int\limits_{0}^{10}[-5-10\cos (\frac{\pi H}{12})]dH\] \[\left[ -5H+\frac{10 \pi}{12}\sin \frac{\pi H}{12} \right]\]Between H=10 and H=0

OpenStudy (kropot72):

The value of the definite integral = -48.7 Therefore the average temperature over the first 10 hours is: -48.7/10 = -4.87 degrees

OpenStudy (anonymous):

sorry, i got off here, but it makes sense using the integration, i knew i had to but wasn't sure how to do it. thanks for helping